Grade 12th passMechanics

A ball of mass 0.1kg makes an elastic head on collision with a ball of unknown mass,initially at rest.If 0.1kg ball rebounds at one third of its original speed what is the mass of other ball?

V G Munj
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
For elastic collision
Let V1 initial = v
V1(final) = (m1-m2)* V/(m1+m2)
Then V1(final) = V/3
V/3 = (0.1-m2) *V/(0.1+m2)
0.1 + m2 = 0.3 - 3 m2
4m2 = 0.2
m2 = 0.05 kg
 
Regards
Arun (askIITians forum expert)